OpenStudy (anonymous):

What is the solution of the system? Use a graph. y = x + 5 y = –5x – 1

OpenStudy (skullpatrol):

The general form of these equations is y = mx + b

OpenStudy (skullpatrol):

y = mx + b y = x + 5 y = –5x – 1

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

these graphs are killing me. i think i am about to fail my first class ever =(

OpenStudy (skullpatrol):

y = mx + b y = x + 5 y = –5x – 1 What is the value for b in each graph?

OpenStudy (anonymous):

5?

OpenStudy (skullpatrol):

Yes, :-D The value of 5 is called the y-intercept of the graph, because it is where the graph of the line intercepts the y-axis.

OpenStudy (skullpatrol):

What is the y-intercept for the second graph? y = –5x – 1

OpenStudy (anonymous):

I'm just going to have to guess on these because I dont see the same thing on the graph as in the written

OpenStudy (skullpatrol):

y = mx + b y = –5x – 1

OpenStudy (skullpatrol):

y = –5x – 1 = -5x + (-1)

OpenStudy (skullpatrol):

y = mx + b y= = -5x + (-1)

OpenStudy (skullpatrol):